OpenStudy (anonymous):

Solve for k. a/b= b/a + 1/k

OpenStudy (anonymous):

a/b=b/a+1/k then (a/b)-(b/a)=1/k then (a^2-b^2)/ab = 1/k then k= (ab)/(a^2-b^2)

OpenStudy (anonymous):

thanks alto

OpenStudy (anonymous):

allot